# On the pro-modularity in the residually reducible case for some totally real fields

### v1

@@ 4.3. The pro-modularity of universal deformation rings
Keep the notations and assumptions in the previous subsection. In this subsection, we further assume that $\bar{\chi}\mid_{G_{F_{v}}}\neq\omega_{p}^{\pm 1}$ for any $v\mid p$, and
$[F:\mathbb{Q}]\geq 7|\Sigma\setminus\Sigma_{p}|+2.$
@@
Lemma 4.3.4.
There exists an irreducible component $C_{1}^{\textnormal{ord}}$ of $C^{\textnormal{ord}}$ such that
1) $\operatorname{dim}C_{1}^{\textnormal{ord}}=[F:\mathbb{Q}]+1$.
2) $C_{1}^{\textnormal{ord}}\to\operatorname{Spec}\Lambda_{F}$ is a surjective map.
3) Let $C_{1}^{\textnormal{ord,aut}}$ be the set of regular de Rham primes in $C_{1}^{\textnormal{ord}}$. Then $C_{1}^{\textnormal{ord,aut}}$ is dense in $C_{1}^{\textnormal{ord}}$. Here a prime $\mathfrak{q}$ is called regular de Rham prime if
Proof.
The proof is similar to [Pan22, Corollary 7.2.3]. We sketch a proof here.
From Proposition 4.2.6, Corollary 4.2.7 and the previous lemma, we can find an irreducible component of dimension $[F:\mathbb{Q}]+1$, and we fix one. As $\Lambda_{F}$ is a domain, the second condition holds.
We claim that this component satisfies the last condition. For $v\mid p$, let $\psi_{v,1}^{\textnormal{ord}},\psi_{v,2}^{\textnormal{ord}}:G_{F_{v}}\to(R^{\textnormal{ps,ord}})^{\times}$ be the liftings of $1,\bar{\chi}\mid_{G_{F_{v}}}$ respectively. By Proposition 4.2.9, there exists $n_{v}\in\{1,2\}$ such that for any $\mathfrak{q}\in C_{1}^{\textnormal{ord}}$, we get $\rho(\mathfrak{q})\mid_{G_{F_{v}}}\cong\begin{pmatrix}\psi_{v,n_{v}}^{\textnormal{ord}}~{}\operatorname{mod}~{}\mathfrak{q}&*\\ 0&\psi_{v,3-n_{v}}^{\textnormal{ord}}~{}\operatorname{mod}~{}\mathfrak{q}\end{pmatrix}$.
Now by Proposition 3.2.3, the subset of $\mathfrak{q}$ such that $\rho(\mathfrak{q})$ is reducible has dimension at most $2+\delta_{F}<1+[F:\mathbb{Q}]$. Hence, we know that the image of $C_{1}^{\textnormal{ord,aut}}$ is dense in $\operatorname{Spec}\Lambda_{F}$, and hence dense in $C_{1}^{\textnormal{ord}}$. ∎
Now let $\mathfrak{Q}$ be the generic point of $C_{1}^{\textnormal{ord}}$. For any $v\in\Sigma\setminus\Sigma_{p}$, we fix a lifting $\sigma_{v}$ of the generator of the $\mathbb{Z}_{p}$-quotient of $I_{v}$ and a lifting $\operatorname{Frob}_{v}$ of the Frobenius element. Write $S=\{\sigma_{v},\operatorname{Frob}_{v}:~{}v\in\Sigma\setminus\Sigma_{p}\}$. Clearly, $|S|=2|\Sigma\setminus\Sigma_{p}|$.
@@
By Proposition 4.2.6, for any prime $\mathfrak{q}_{1}\in C_{1}^{\textnormal{ord,aut}}$, we know that $\rho(\mathfrak{q}_{1})$ comes from a twist of Hilbert modular form. By the density result in Lemma 4.3.4, the image of any prime of $C_{1}^{\textnormal{ord}}$ in $\operatorname{Spec}R^{\textnormal{ps},\{\xi_{v}\}}$ is pro-modular.
@@
(1) By the proof of [Pan22, Lemma 7.1.2], for a one-dimensional prime $\mathfrak{p}$ of $R_{\textnormal{aux}}^{\textnormal{ps}}$, if it corresponds to an irreducible pseudo-representation, then every irreducible component of $R_{\textnormal{aux}}^{\textnormal{ps}}$ containing $\mathfrak{p}$ is of dimension at least $1+2[F:\mathbb{Q}]$. These irreducible components are what we are concerned about in Theorem 4.3.2.
(2) In the case $\bar{\chi}=\omega_{p}^{\pm 1}$, we can also prove a pro-modularity result as Theorem 4.3.2, but the deduction is more complicated. Our method seems to be helpful to prove the two-dimensional Fontaine-Mazur conjecture in the regular case when $p=3$ and $\bar{\chi}=\omega_{3}$. See [Zha24] for more details.
@@
then for any non-zero element $x\in H^{1}(F_{\Sigma}/F,\mathbb{F}(\bar{\chi}^{-1}))$, any prime ideal of $R_{x}^{\textnormal{aux}}$ is pro-modular.
$(2)$ Keep the assumption in $(1)$. If the natural surjection $R^{\textnormal{ps},\{\xi_{v}\}}\twoheadrightarrow(\mathbb{T}_{\xi_{v}})_{\mathfrak{m}_{\xi_{v}}}$ factors through $R_{x}^{\textnormal{aux},\{\xi_{v}\}}$ for some non-zero cocycle $x\in H^{1}(F_{\Sigma}/F,\mathbb{F}(\bar{\chi}^{-1}))$, then we have
@@
Let $\mathfrak{P}$ be a minimal prime of $R_{x}^{\textnormal{aux}}$. Let $\mathfrak{P}^{\prime}$ be the inverse image of $\mathfrak{P}$ via the natural map $f:R_{\textnormal{aux}}^{\textnormal{ps}}\to R_{x}^{\textnormal{aux}}$. From our assumptions, Proposition 3.1.3, Proposition 3.1.4 and Proposition 3.2.5, we know that $\mathfrak{P}^{\prime}$ is a minimal prime of dimension at least $1+2[F:\mathbb{Q}]$. By Theorem 4.3.2, we know that $\mathfrak{P}$ is pro-modular. This proves (1).
@@
5. Applications
In this section, we use our pro-modularity arguments to study the dimensions of the big Hecke algebra and Galois deformation rings.
5.1. The dimension of Hecke algebra
In this subsection, we study the dimension of our big Hecke algebra $(\mathbb{T}_{\xi})_{\mathfrak{m}_{\xi}}$.
In Deo’s article [Deo23], it is proved that the big Hecke algebra is a reduced local complete intersection ring of Krull dimension $4$. One of the crucial step of his proof is the following infinite fern argument.
@@ 5.1. The dimension of Hecke algebra
[Eme11, Corollary 2.28] The Krull dimension of each component of the spectrum of the big Hecke algebra for $\mathbb{Q}$ is at least $4$.
@@
Such a result is first studied by Gouvêa and Mazur in [GM97]. The case for totally real fields is studied in [Che11].
Using Pan’s patching theorem and our pro-modularity arguments, we can give an analogue result to Deo’s work.
Theorem 5.1.3.
Keep assumptions as in Theorem 4.3.2. Then $(\mathbb{T}_{\xi})_{\mathfrak{m}_{\xi}}$ is equidimensional of dimension $1+2[F:\mathbb{Q}]$.
Before giving the proof, we briefly summarize some main steps of the proof of Theorem 4.2.4. Here we follow most notations as in [Pan22, Section 4], and one can find more details there.
Using Taylor’s Ihara avoidance trick [Tay08], one can show that $\widehat{(R_{\infty}^{\{\xi_{v}\}})_{\mathfrak{q}_{\infty}^{\{\xi_{v}\}}}}$ is equidimensional of dimension $4|P|-1+r+2[F:\mathbb{Q}]$ (see [Pan22, Section 4.2, Lemma 4.8.2]). As $(R^{\{\xi_{v}\}})^{\prime}$ is finite free over $\widehat{(R_{\infty}^{\{\xi_{v}\}})_{\mathfrak{q}_{\infty}^{\{\xi_{v}\}}}}$, we have
From Theorem 4.2.1, [Eis13, Corollary 12.5], and discussions above, we can also conclude that for a nice prime $\mathfrak{q}$ of $(\mathbb{T}_{\xi})_{\mathfrak{m}_{\xi}}$, we have $\operatorname{dim}\widehat{(\mathbb{T}_{\xi})_{\mathfrak{q}}}=\operatorname{dim}(\mathbb{T}_{\xi})_{\mathfrak{q}}=2[F:\mathbb{Q}]$.
Proof of Theorem 5.1.3.
Similar to [Pan22, Corollary 3.3.3], we know that $(\mathbb{T}_{\xi})_{\mathfrak{m}_{\xi}}$ is a CNL ring and hence catenary. Then for every irreducible component of $(\mathbb{T}_{\xi})_{\mathfrak{m}_{\xi}}$, if it contains a nice prime, it is of dimension $1+2[F:\mathbb{Q}]$.
From the proof of Theorem 4.3.2, we know that every irreducible component of $R^{\textnormal{ps},\{\xi_{v}\}}$ of dimension at least $1+2[F:\mathbb{Q}]$ contains a nice prime. Combining Theorem 4.2.1 and Theorem 4.2.4, we know that for every irreducible component of $(\mathbb{T}_{\xi})_{\mathfrak{m}_{\xi}}$ contains a nice prime. This proves our result. ∎
5.2. The dimension of Galois deformation ring
In this subsection, we study the dimension of $R_{x}^{\textnormal{aux}}$.
Keep assumptions as in Theorem 4.3.2. The following result is an analogue to [Deo23, Theorem A] and Proposition 3.1.5.
Theorem 5.2.1.
$(1)$ For every irreducible component of $R_{\textnormal{aux}}^{\textnormal{ps}}$ of dimension at least $1+2[F:\mathbb{Q}]$, it is of dimension $1+2[F:\mathbb{Q}]$.
$(2)$ As in Theorem 4.3.7, we further assume
$[F:\mathbb{Q}]\geq\operatorname{max}\{1+\frac{1}{2}\operatorname{dim}_{\mathbb{F}}(H^{1}(F_{\Sigma}/F,\mathbb{F}(\bar{\chi}^{-1}))),7|\Sigma\setminus\Sigma_{p}|+2\}.$
Then for any non-zero element $x\in H^{1}(F_{\Sigma}/F,\mathbb{F}(\bar{\chi}^{-1}))$, $R_{x}^{\textnormal{aux}}$ is a local complete intersection ring of Krull dimension $1+2[F:\mathbb{Q}]$.
Proof.
The first assertion follows from Theorem 4.2.4 and Theorem 5.1.3.
Let $\mathfrak{P}$ be a minimal prime of $R_{x}^{\textnormal{aux}}$. Let $\mathfrak{P}^{\prime}$ be the inverse image of $\mathfrak{P}$ via the natural map $f:R_{\textnormal{aux}}^{\textnormal{ps}}\to R_{x}^{\textnormal{aux}}$. From our assumptions, Proposition 3.1.3, Proposition 3.1.4 and Proposition 3.2.5, we have
$1+2[F:\mathbb{Q}]\leq\operatorname{dim}R_{x}^{\textnormal{aux}}/\mathfrak{P}\leq\operatorname{dim}R_{\textnormal{aux}}^{\textnormal{ps}}/\mathfrak{P}^{\prime}.$
Then by the first assertion, we know that $\mathfrak{P}^{\prime}$ is of dimension $1+2[F:\mathbb{Q}]$. Consequently, $R_{x}^{\textnormal{aux}}$ is of dimension at most $1+2[F:\mathbb{Q}]$. Using Proposition 3.1.2 and Proposition 3.1.3, we obtain that $R_{x}^{\textnormal{aux}}$ is a local complete intersection ring of Krull dimension $1+2[F:\mathbb{Q}]$.
Acknowledgement.
Appendix: Galois cohomology
Let $\theta:\operatorname{Gal}(F_{\Sigma}/F)\to\mathbb{F}^{\times}$ be a continuous totally odd character. Write $F_{\theta}$ be the splitting field of the character $\theta\omega_{p}^{-1}$. Let $\operatorname{Cl}(F_{\theta})$ be the $p$-class group of $F_{\theta}$.
Proposition.
We have $\operatorname{dim}_{\mathbb{F}}(H^{1}(F_{\Sigma}/F,\mathbb{F}(\theta^{-1})))=\operatorname{dim}_{\mathbb{F}}(H^{1}(F_{\Sigma}/F,\mathbb{F}(\theta)))=[F:\mathbb{Q}]$ if and only if the following condition holds:
1) $\theta\mid_{G_{F_{v}}}\neq\omega_{p}^{\pm 1}\mid_{G_{F_{v}}}$ for all $v\in\Sigma$,
2) $\operatorname{Cl}(F_{\theta})[\theta\omega_{p}^{-1}]=\operatorname{Cl}(F_{\theta})[\theta^{-1}\omega_{p}]=0$, where $\operatorname{Cl}(F_{\theta})[\eta]$ is the $\eta$-eigenspace of $\operatorname{Cl}(F_{\theta})$ for any character $\eta:\operatorname{Gal}(F_{\theta}/F)\to\mathbb{F}^{\times}$.
Proof.
It is an analogue to [Deo23, Lemma 2.21]. We sketch a proof here.
Let $H_{\Sigma}=\operatorname{ker}(H^{1}(F_{\Sigma}/F,\mathbb{F}(\theta^{-1}\omega_{p}))\to\prod_{v\in\Sigma}H^{1}(F_{v},\mathbb{F}(\theta^{-1}\omega_{p})))$ be the ”Selmer” group. By the Greenberg-Wiles formula ([NSW13, Theorem 8.7.9]) and local characteristic formula, we have
$\operatorname{dim}_{\mathbb{F}}(H^{1}(F_{\Sigma}/F,\mathbb{F}(\theta)))=\operatorname{dim}_{\mathbb{F}}(H_{\Sigma})+[F:\mathbb{Q}]+\sum_{v\in\Sigma}\operatorname{dim}_{\mathbb{F}}(H^{0}(F_{v},\mathbb{F}(\theta^{-1}\omega_{p}))),$
where we take the local condition $\mathcal{L}_{v}$ as $H^{1}(F_{v},\theta\mid_{G_{F_{v}}})$. Thus, we have $\operatorname{dim}_{\mathbb{F}}(H^{1}(F_{\Sigma}/F,\mathbb{F}(\theta)))=[F:\mathbb{Q}]$ if and only if the following conditions hold:
i) $H^{0}(F_{v},\mathbb{F}(\theta^{-1}\omega_{p}))=0$ for all $v\in\Sigma$,
ii) the Selmer group $H_{\Sigma}$ is trivial.
Clearly, i) is equivalent to $\theta\mid_{G_{F_{v}}}\neq\omega_{p}\mid_{G_{F_{v}}}$ for all $v\in\Sigma$.
By inflation-restriction exact sequence, we have
$H^{1}(F_{v}^{ur}/F_{v},(\mathbb{F}(\theta^{-1}\omega_{p}))^{I_{v}})=\operatorname{ker}(H^{1}(F_{v},\mathbb{F}(\theta^{-1}\omega_{p}))\to H^{1}(I_{v},\mathbb{F}(\theta^{-1}\omega_{p}))),~{}v\in\Sigma,$
where $F_{v}^{ur}$ is the maximal unramified extension of $F_{v}$. As $\mathbb{F}$ is a finite field, we have
$(\mathbb{F}(\theta^{-1}\omega_{p}))^{I_{v}}/(\operatorname{Frob}_{v}-1)(\mathbb{F}(\theta^{-1}\omega_{p}))^{I_{v}}=H^{1}(F_{v}^{ur}/F_{v},(\mathbb{F}(\theta^{-1}\omega_{p}))^{I_{v}})=0,$
if $\theta\mid_{G_{F_{v}}}\neq\omega_{p}\mid_{G_{F_{v}}}$. Therefore,
$H_{\Sigma}=\operatorname{ker}(H^{1}(F_{\Sigma}/F,\mathbb{F}(\theta^{-1}\omega_{p}))\to\prod_{v\in\Sigma}H^{1}(I_{v},\mathbb{F}(\theta^{-1}\omega_{p}))).$
By class field theory, we know that $\operatorname{dim}_{\mathbb{F}}(H^{1}(F_{\Sigma}/F,\mathbb{F}(\theta)))=[F:\mathbb{Q}]$ if and only if the following conditions hold:
i’) $\theta\mid_{G_{F_{v}}}\neq\omega_{p}\mid_{G_{F_{v}}}$ for all $v\in\Sigma$,
ii’) $\operatorname{Cl}(F_{\theta})[\theta^{-1}\omega_{p}]=0$.
Our result follows from this equivalence. ∎

### v3

@@ 4.4. The pro-modularity of universal deformation rings
Keep the notations and assumptions on the finite set $\Sigma$ and the character $\chi$ in the previous two subsections. In this subsection, we further assume that $\bar{\chi}\mid_{G_{F_{v}}}\neq\omega_{p}^{\pm 1}$ for any $v\mid p$, and $[F:\mathbb{Q}]\geq 7|\Sigma\setminus\Sigma_{p}|+2.$
@@
Now we can choose an irreducible component $C_{1}^{\textnormal{ord}}\subset C^{\textnormal{ord}}$ of $\operatorname{Spec}R^{\textnormal{ps,ord}}$ of dimension $1+[F:\mathbb{Q}]$, and we let $\mathfrak{Q}$ be the generic point of $C_{1}^{\textnormal{ord}}$. For any $v\in\Sigma\setminus\Sigma_{p}$, we fix a lifting $\sigma_{v}$ of the generator of the $\mathbb{Z}_{p}$-quotient of $I_{v}$ and a lifting $\operatorname{Frob}_{v}$ of the Frobenius element. Write $S=\{\sigma_{v},\operatorname{Frob}_{v}:~{}v\in\Sigma\setminus\Sigma_{p}\}$. Clearly, $|S|=2|\Sigma\setminus\Sigma_{p}|$.
@@
By Proposition 4.3.3, for any prime $\mathfrak{q}_{1}\in C_{1}^{\textnormal{ord,aut}}$, the subset of $C_{1}^{\textnormal{ord}}$ consisting of all regular de Rham primes as defined in Lemma 4.3.4, we know that $\rho(\mathfrak{q}_{1})$ comes from a twist of Hilbert modular form. By the density result in Lemma 4.3.4, the image of any prime of $C_{1}^{\textnormal{ord}}$ in $\operatorname{Spec}R^{\textnormal{ps},\{\xi_{v}\}}$ is pro-modular.
@@
1) By Proposition 3.2.7, for a one-dimensional prime $\mathfrak{p}$ of $R_{\textnormal{aux}}^{\textnormal{ps}}$, if it corresponds to an irreducible pseudo-representation, then at least one of the irreducible components of $R_{\textnormal{aux}}^{\textnormal{ps}}$ containing $\mathfrak{p}$ is of dimension at least $1+2[F:\mathbb{Q}]$. Hence $\mathfrak{p}$ is pro-modular.
2) In the proof of Theorem 4.4.2, we can also find that the irreducible component $C$ is of dimension $1+2[F:\mathbb{Q}]$. Note that the Iwasawa algebra $\Lambda_{F}$ is of dimension $1+[F:\mathbb{Q}]$. By the finiteness result (Proposition 4.3.3), we have $\operatorname{dim}C^{\textnormal{ord}}\leq[F:\mathbb{Q}]+1$. Combining Lemma 4.4.3, we have $\operatorname{dim}C^{\textnormal{ord}}=[F:\mathbb{Q}]+1$, and hence $\operatorname{dim}C=2[F:\mathbb{Q}]+1$.
@@
then for any non-zero element $x\in H^{1}(F_{\Sigma}/F,\mathbb{F}(\bar{\chi}^{-1}))$, any prime ideal of $R_{x}^{\textnormal{aux}}$ is pro-modular. Moreover, $R_{x}^{\textnormal{aux}}$ is a local complete intersection ring of Krull dimension $1+2[F:\mathbb{Q}]$.
$(2)$ Keep the assumption as in $(1)$. If the natural surjection $R^{\textnormal{ps},\{\xi_{v}\}}\twoheadrightarrow(\mathbb{T}_{\xi_{v}})_{\mathfrak{m}_{\xi_{v}}}$ factors through $R_{x}^{\textnormal{aux},\{\xi_{v}\}}$ for some non-zero cocycle $x\in H^{1}(F_{\Sigma}/F,\mathbb{F}(\bar{\chi}^{-1}))$, then we have
@@
Let $\mathfrak{P}$ be a minimal prime of $R_{x}^{\textnormal{aux}}$. Let $\mathfrak{P}^{\prime}$ be the inverse image of $\mathfrak{P}$ via the natural map $f:R_{\textnormal{aux}}^{\textnormal{ps}}\to R_{x}^{\textnormal{aux}}$. From our assumptions, Proposition 3.1.3, Proposition 3.1.4 and Proposition 3.2.5, we know that $\mathfrak{P}^{\prime}$ is a minimal prime of dimension at least $1+2[F:\mathbb{Q}]$. By Theorem 4.4.2, we know that $\mathfrak{P}$ is pro-modular.
By Remark 4.4.5, we know that $\mathfrak{P}^{\prime}$ is of dimension $1+2[F:\mathbb{Q}]$. Consequently, $R_{x}^{\textnormal{aux}}$ is of dimension at most $1+2[F:\mathbb{Q}]$. Combining Proposition 3.1.2 and Proposition 3.1.3, we obtain that $R_{x}^{\textnormal{aux}}$ is a local complete intersection ring of Krull dimension $1+2[F:\mathbb{Q}]$. This proves (1).
@@
4.5. A special case
In this subsection, we deal with the case $\bar{\chi}\mid_{G_{F_{v}}}=\omega_{p}$ for any $v|p$. In this case, the deduction is more complicated because Lemma 4.4.3 no longer holds. We follow the strategy of the proof in [Pan22, Section 7.4].
Keep assumptions on as in Section 4.2 and Section 4.3 the finite set $\Sigma$ and the character $\chi$ in the previous two subsections. We further assume $[F:\mathbb{Q}]\geq 7|\Sigma\setminus\Sigma_{p}|+4.$
Proposition 4.5.1.
Let $\mathfrak{r}$ be a one-dimensional irreducible pro-modular prime of $R_{\textnormal{aux}}^{\textnormal{ps}}$. Then every irreducible component of $R_{\textnormal{aux}}^{\textnormal{ps}}$ containing $\mathfrak{r}$ contains a nice prime, and hence is pro-modular.
Proof.
Let $Z$ be the set of all irreducible components of $(R_{\textnormal{aux}}^{\textnormal{ps}})_{\mathfrak{r}}$. Write $Z=Z_{1}\amalg Z_{2}$, where $Z_{1}$ is the subset of $Z$ consisting of all pro-modular irreducible components. We first show that if $Z_{1}$ is not empty, then $Z_{2}$ is empty.
If not, by Proposition 3.2.7, there exist an irreducible component $Y\in Z_{2}$ and a pro-modular prime $\mathfrak{r}^{\prime}\in Y$ of dimension at least $2[F:\mathbb{Q}]-1$.
Arguing as the proof of Theorem 4.4.2, we may find a prime $\mathfrak{r}^{\prime\prime}\subset\mathfrak{r}^{\prime}$ satisfying the following conditions:
  • $\varpi\in\mathfrak{r}^{\prime\prime}$.
  • The dimension of $\mathfrak{r}^{\prime\prime}$ is at least $2[F:\mathbb{Q}]-7|\Sigma\setminus\Sigma_{p}|-2\geq[F:\mathbb{Q}]+2$.
  • For any irreducible one-dimensional prime $\mathfrak{t}$ containing $\mathfrak{r}^{\prime\prime}$, there exists a $2$-dimensional Galois representation $\rho(\mathfrak{t}):G_{F,\Sigma}\to\operatorname{GL}_{2}(k(\mathfrak{t}))$ satisfying the conditions in (2) of Definition 4.2.3 except the third one.
If $\bar{\chi}$ is not quadratic, then by Lemma 4.2.4, we can find a nice prime containing $\mathfrak{r}^{\prime\prime}$ (hence containing $\mathfrak{r}^{\prime}$). Using Theorem 4.2.5, we conclude that $Y$ is pro-modular, which contradicts to our definition of $Z_{2}$.
If $\bar{\chi}$ is quadratic, we know that the splitting field $F(\bar{\chi})$ is an abelian CM field. Then by Corollary 3.2.3, we may find a prime containing $\mathfrak{r}^{\prime}$, whose corresponding Galois representation is irreducible and not dihedral. From our choice of $\mathfrak{r}^{\prime\prime}$, we know that this is actually a nice prime. Using Theorem 4.2.5 again, we get a contradiction.
As $\mathfrak{r}$ is pro-modular, using Corollary 4.2.1, we know that there exists a pro-modular prime $\mathfrak{r}_{1}$ contained in $\mathfrak{r}$ of dimension at least $1+2[F:\mathbb{Q}]$. Arguing as above, we can find a nice prime containing $\mathfrak{r}_{1}$. This implies that at least one of the irreducible component containing $\mathfrak{r}$ is pro-modular. Hence, $Z_{1}$ is not empty. ∎
From now on, we assume $\bar{\chi}\mid_{G_{F_{v}}}=\omega_{p}$ for any $v|p$. Suppose that $C$ is an irreducible component of $R_{\textnormal{aux}}^{\textnormal{ps}}$ of dimension at least $1+2[F:\mathbb{Q}]$. In this case, by [Pan22, Lemma 7.4.4], the kernel of the natural surjection $R_{v}^{\textnormal{ps}}\twoheadrightarrow R_{v}^{\textnormal{ps,ord}}$ can be generated by two elements. Then by Krull’s principal ideal theorem, we have $\operatorname{dim}C^{\textnormal{ord}}\geq 1$.
Choose a prime $\mathfrak{q}\in C^{\textnormal{ord}}$ such that $\dim R_{\textnormal{aux}}^{\textnormal{ps}}/\mathfrak{q}=1$. Enlarging $\mathcal{O}$ if necessary, we may assume the normalization of $R_{\textnormal{aux}}^{\textnormal{ps}}/\mathfrak{q}$ is either $\mathcal{O}$ or isomorphic to $\mathbb{F}[[T]]$. There are three possibilities for the associated semi-simple representation $\rho(\mathfrak{q})$:
  1. (1)
    $\rho(\mathfrak{q})$ is irreducible.
  2. (2)
    $\rho(\mathfrak{q})\cong\psi_{1}\oplus\psi_{2}$ is reducible and $\psi_{1}/\psi_{2}$ is not of the form $\epsilon^{\pm 1}\theta$, where $\theta$ is a finite order character of $G_{F}$. We call this case generic reducible.
  3. (3)
    $\rho(\mathfrak{q})\cong\psi_{1}\oplus\psi_{2}$ and $\psi_{1}/\psi_{2}=\varepsilon\theta$, for some finite order character $\theta$ of $G_{F}$. We call this case non-generic reducible.
@@ 4.5. A special case
The irreducible component $C$ is pro-modular of dimension $1+2[F:\mathbb{Q}]$.
Proof.
As the discussions above, we divide the proof into three cases.
1. The irreducible case
Using the same proof of [Pan22, Lemma 7.4.7], we find that the localization $(R^{\textnormal{ps,ord}})_{\mathfrak{q}}$ is of dimension at least $[F:\mathbb{Q}]$. In other words, there exists an irreducible component of $R^{\textnormal{ps,ord}}$ containing $\mathfrak{q}$ of dimension at least $1+[F:\mathbb{Q}]$. Arguing as in the proof of Theorem 4.4.2, we know that $\mathfrak{q}$ is pro-modular. By Proposition 4.5.1 and Remark 4.2.6, we know that $C$ is pro-modular of dimension $1+2[F:\mathbb{Q}]$.
2. The generic reducible case
By the same proof as in [Pan22, Section 7.4.4], we have $\operatorname{dim}C^{\textnormal{ord}}\geq 1+[F:\mathbb{Q}]$ (see [Pan22, Corollary 7.4.9]). Thus, as in the irreducible case, we also have $C$ is pro-modular of dimension $1+2[F:\mathbb{Q}]$.
3. The non-generic reducible case
We only sketch the proof.
For any non-zero class $B\in\operatorname{Ext}^{1}_{E[G_{F,\Sigma}]}(\psi_{1},\psi_{2})\cong H^{1}(G_{F,\Sigma},E(\psi_{2}/\psi_{1}))$, it corresponds to a non-split extension $\rho_{B}:G_{F,\Sigma}\to\operatorname{GL}_{2}(E)$ of $\psi_{1}$ by $\psi_{2}$. We may assume $\rho_{B}$ is of the form
$\begin{pmatrix}\psi_{2}&b_{B}\\ 0&\psi_{1}\end{pmatrix}.$
We denote the universal deformation ring of $\rho_{B}$ with fixed determinant $\chi$ by $R_{B}$.
We define $Z_{B}$ as the set of irreducible components of $R_{\textnormal{aux}}^{\textnormal{ps}}$ whose generic points lie in the image of $\operatorname{Spec}R_{B}\to\operatorname{Spec}R_{\textnormal{aux}}^{\textnormal{ps}}$.
First, by the arguments in [Pan22, Step (1), page 1153], there exists a non-zero element $B_{0}$ in $H^{1}(G_{F,\Sigma},E(\psi_{2}/\psi_{1}))$ such that we can find a pro-modular irreducible component in $Z_{B_{0}}$. Using [Pan22, Corollary 7.4.13] and Proposition 4.5.1, we know that any irreducible component in $Z_{B_{0}}$ is pro-modular.
Second, using the same proof in [Pan22, Step (2), page 1154 & 1155], we can find an element $B_{1}$ in $H^{1}(G_{F,\Sigma},E(\psi_{2}/\psi_{1}))$ such that $C\in Z_{B_{1}}$ and there exist irreducible components $C_{B_{0}}\in Z_{B_{0}},C_{B_{1}}\in Z_{B_{1}}$ containing a prime of dimension $\mathfrak{s}$ at least $2+[F:\mathbb{Q}]$. As $C_{B_{0}}$ is pro-modular, we know that $\mathfrak{s}$ is pro-modular. By Corollary 3.2.3, we can find a one-dimensional irreducible pro-modular prime in $C_{B_{0}}\cap C_{B_{1}}$. Combining Proposition 4.5.1, we know that $C_{B_{1}}$ is pro-modular. Using [Pan22, Corollary 7.4.13] and Proposition 4.5.1 again, we know that any irreducible component in $Z_{B_{1}}$ is pro-modular. In particular, $C$ is pro-modular of dimension $1+2[F:\mathbb{Q}]$. ∎
@@ 3. The non-generic reducible case
1) If $\bar{\chi}$ is quadratic, then we cannot ignore the dihedral condition in the defintion of a nice prime. By Corollary 3.2.2, the upper bound of the dihedral locus is $1+[F:\mathbb{Q}]$, so we may not find a nice prime from the pro-modular prime $\mathfrak{s}$ as its dimension may not be large enough. However, we can find one-dimensional irreducible pro-modular prime from $\mathfrak{s}$ and then use Proposition 4.5.1, which is the advantage of our method (particularly Corollary 2.1.7) in this paper.
2) We can also prove the same results as Theorem 4.4.6 using the same proof.